1. A company is deciding whether or not to buy a machine. The machine costs $45,000 and is expected to generate net cash flow for the business as follows: Year 1 $12,000 Year 2 $18,000 Year 3 $26,000 The company’s applicable interest rate is 12% on the machine That is, the company will only invest in the machine if the cash flow it receives generates a return of 12% or more. (In practice this rate is often based on the company’s cost of capital). Should the company buy the machine? Why or why not?
Answer #1

Present value of cash inflow = (12000*0.89286+18000*0.79719+26000*0.71178) = $43570.02

Net present value = Present value of cash inflow-Present value of cash outflow

= 43570.02-45000

Net present value = -1429.98

No, Company should not buy the machine because net present value is negative. or in other words present value of cash inflow is lower than present value of cash outflow